# Using Diffusion Models to do Data Assimilation

### v1

@@ 3.2.1 Bayesian posterior
For our simplified linear example in which the observation network and observation error variance is fixed in time, it is well-known that the forecast covariance, analysis (posterior) covariance and Kalman gain converge to a steady-state, i.e., $\mathbf{P}^{f}\to\mathbf{P}^{f}_{\infty}$, $\mathbf{P}_{a}\to\mathbf{P}^{a}_{\infty}$, $\mathbf{K}\to\mathbf{K}_{\infty}$ and therefore we have that
@@
$\mathbf{K}_{\infty}$ $=\mathbf{P}^{f}_{\infty}\mathbf{H}^{T}\left(\mathbf{H}\mathbf{P}^{f}_{\infty}\mathbf{H}^{T}+r\mathbf{I}\right)^{-1}.$ (41)
The posterior mean at time $k$ is
@@
where $\overline{\mathbf{x}}_{fk}$ is the forecast mean, i.e., the posterior mean of the previous cycle evolved forward using an ensemble under the dynamics (17) to the next observation time. Note that our focus on the steady-state here can be understood as having completed sufficiently many cycles with a well-constructed conventional DA system.
Since our observation system is simple (direct observations of $n_{y}$ elements of $\mathbf{x}_{k}$), it is again sufficient to focus on one observed variable from the state vector. Denoting the $j^{th}$ diagonal element of $\mathbf{P}^{f}_{\infty}$ by $\alpha_{j}$ we obtain an expression for the $j^{th}$ element of the analysis mean at time $k$
$[\overline{\mathbf{x}}^{a}_{k}]^{j}=[\overline{\mathbf{x}}_{fk}]^{j}+\frac{\alpha^{j}}{\alpha^{j}+r}([y_{k}]^{j}-[\overline{\mathbf{x}}_{fk}]^{j}).$ (43)
The $j^{th}$ diagonal element of the posterior covariance becomes
@@ 3.2.2 Diffusion DA
Here we will extend the typical training paradigm of diffusion modeling to the Bayesian posterior distribution with a cycling prior, i.e., the posterior distribution typically targeted in conventional DA. As explained before, the “training set” for ML is typically taken to be the climatological prior, but in the simplified problem setup we are working in, one can imagine a diffusion DA system that re-trains at every cycle (see Bao et al. (2024) for another example).
@@
    We use this diffusion model to generate a large training set consistent with $p(\mathbf{x}_{0}|\mathbf{y}_{0})$.
@@
    We use the forecast model (17) to push each member of this set forward to the time of the next set of observations, $\mathbf{y}_{1}$.
@@
    We then use these forecasts as the training set to build a new diffusion model that produces samples from $p(\mathbf{x}_{1}|\mathbf{y}_{1},\mathbf{y}_{0})$.
@@
    We then repeat steps 3 and 4 for each cycle $k$ for which we desire to process observations.
We imagine we have done this up to the $k^{th}$ cycle and that the covariances have converged to their steady-state values. In this case, the typical diffusion model loss function would be modified to draw its training set from the $k^{th}$ cycling joint posterior, i.e.,
$p(\mathbf{x}_{k},\mathbf{y}_{k}|\mathbf{y}_{k-1},\mathbf{y}_{k-2},...)=p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k}|\mathbf{y}_{k-1},\mathbf{y}_{k-2},...)$ (45)
where we emphasize that $p(\mathbf{x}_{k}|\mathbf{y}_{k-1},\mathbf{y}_{k-2},...)$ is the cycling prior and $p(\mathbf{y}_{k}|\mathbf{x}_{k})$ is the standard Gaussian observation likelihood. This implies the following loss
$\ell=E_{\mathbf{x}_{k},\mathbf{y}_{k}\sim p(\mathbf{x}_{k},\mathbf{y}_{k}|\mathbf{y}_{k-1},\mathbf{y}_{k-2},...)}E_{n\sim\mathcal{N}(0,t^{2})}||\mathbf{x}_{k}-\mathbf{D}(\mathbf{v}=\mathbf{x}_{k}+n,\mathbf{y}_{k},t)||_{2}^{2},$ (46)
Note that this loss function implies that we re-train the diffusion model at each and every cycle.
In any event, the minimum of this loss function for this linear, Gaussian system is simply
$\mathbf{D}(\mathbf{v},\mathbf{y}_{k};t)=E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{y}_{k},\mathbf{y}_{k-1},...)}[\mathbf{x}_{k}]=\overline{\mathbf{x}}_{fk}+\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}$ (47)
where
$\mathbf{R}=\begin{bmatrix}t^{2}\mathbf{I}_{n_{x}}&\mathbf{0}\\ \mathbf{0}&r\mathbf{I}_{n_{o}}\end{bmatrix},\quad\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}-\overline{\mathbf{x}}_{fk}\\ \mathbf{y}-\mathbf{H}\overline{\mathbf{x}}_{fk}\end{bmatrix}.$ (48)
We then apply Tweedie’s formula to find the following SDE:
$\text{d}\mathbf{v}=-\frac{2}{t}\left[\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}-(\mathbf{v}-\overline{\mathbf{x}}_{fk})\right]\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t},$ (49)
As we did in the previous sections we consider the $j^{th}$ observed grid point. In this case, the reverse process can be written as
@@
We first compute the mean:
$\langle v(0)\rangle=\overline{x}+\frac{\frac{\alpha^{j}r}{\alpha^{j}+r}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2}}(\langle v(T)\rangle-\overline{x})+\frac{(\alpha^{j})^{2}r}{(\alpha^{j}+r)^{2}}\frac{T^{2}}{\frac{\alpha^{j}r}{\alpha^{j}+r}(\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2})}y.$ (52)
In the limit as $T\rightarrow\infty$ the mean simplifies to
$\langle v(0)\rangle_{T\rightarrow\infty}=\overline{x}+\frac{\alpha^{j}}{\alpha^{j}+r}(y-\overline{x})$ (53)
which is equal to the cycling posterior mean derived via the Kalman filter in (43).
Second, we find the variance,
@@
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{\frac{\alpha^{j}r}{\alpha^{j}+r}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2}}\langle(v(T)-\langle v(T)\rangle)^{2}\rangle+\Bigg{\langle}\Bigg{(}\frac{\alpha^{j}r}{\alpha^{j}+r}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\end{split}$ (54)
where we have used the fact that the cross-term vanishes. The stochastic integral on the right-hand side requires the use $\hat{\textrm{o}}$ isometry, i.e.,
$\begin{split}\Bigg{\langle}\Bigg{(}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}=\Bigg{\langle}\Bigg{(}-\int_{0}^{T}\frac{\sqrt{2t}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\\ =2\int_{0}^{T}\frac{t}{(\frac{\alpha^{j}r}{\alpha^{j}+r}+t^{2})^{2}}\text{d}t=-2\int_{T}^{0}\frac{t}{(\frac{\alpha^{j}r}{\alpha^{j}+r}+t^{2})^{2}}\text{d}t\end{split}$ (55)
Thus, the variance becomes
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{\frac{(\alpha^{j})^{2}r^{2}}{(\alpha^{j}+r)^{2}}T^{2}}{(\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2})^{2}}+\frac{\alpha^{j}r}{\alpha^{j}+r}\frac{T^{2}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2}}\end{split}.$ (56)
In the limit as $T\rightarrow\infty$ the variance is
$\langle(v(0)-\langle v(0)\rangle)^{2}\rangle_{T\rightarrow\infty}=\frac{\alpha^{j}r}{\alpha^{j}+r}$ (57)
which is the posterior variance we obtained via the Kalman filter in (44). We have thus shown that retraining a diffusion model at each DA cycle, using the latest prior, results in a diffusion DA system that samples the same posterior pdf as a conventional ensemble DA system that uses a cycling prior.
3.3 DA with an extended likelihood
We now consider DA systems that ingest the observations as well as a forecast, but treat this forecast essentially as an additional observation. We will denote the forecast at time $k$ by $\mathbf{f}_{k}$ and assume that it is derived from a single model forecast from the posterior mean at time $k-1$. In this way we will bring information from the past into the latest state estimate.
If we assume the usual observation equation (18), we have that $p(\mathbf{y}_{k}|\mathbf{x}_{k},\mathbf{f}_{k})=p(\mathbf{y}_{k}|\mathbf{x}_{k})$ so that (15) becomes
$p(\mathbf{x}_{k}|\mathbf{y}_{k},\mathbf{f}_{k})\propto p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{f}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k}).$ (58)
Here, $p(\mathbf{x}_{k})$ is the climatological prior, $p(\mathbf{y}_{k}|\mathbf{x}_{k})$ is the usual Gaussian observation likelihood and $p(\mathbf{f}_{k}|\mathbf{x}_{k})$ is the extension of the likelihood to include the forecast $\mathbf{f}_{k}$.
While it is clear what the climatological prior and Gaussian observation likelihood are, at first blush it is far less clear what the forecast likelihood is. Note however that in this linear, Gaussian system the posterior mean at time $k-1$ is a linear function of the observation (see equation (23)), which, because the observation is a linear function of the truth (see equation (18)), implies that the forecast is also a linear function of the truth. Furthermore, for the example problem of this section, the Kalman gain is a number less than one and the observation is multiplied by this Kalman gain (again see equation (23)). Hence, there will be a linear relation between the forecast $\mathbf{f}_{k}$ and the “truth” $\mathbf{x}_{k}$ of the form
$\mathbf{f}_{k}=a\mathbf{x}_{k}+\boldsymbol{\varepsilon}_{f},\quad\boldsymbol{\varepsilon}_{f}\sim\mathcal{N}(\mathbf{0},r_{f}\mathbf{I}),$ (59)
where $a\leq 1$ and $r_{f}$ are scalars. Note that $r_{f}$ denotes the error variance of the forecast and $a$ is determined from the steady-state Kalman gain, $\mathbf{K}_{\infty}$, and the drift term in (17). In the numerical illustration below we will carefully discuss how we determined the scalars $a$ and $r_{f}$. Finally, we emphasize that this form for $p(\mathbf{f}_{k}|\mathbf{x}_{k})$ is entirely dependent on the linear, Gaussian nature of (17). Consequently, in a nonlinear problem the structure of (59) would generally require a nonlinear function of $\mathbf{x}_{k}$ and a non-Gaussian error distribution. This would be very difficult to determine explicitly in a typical geophysical system, but could be learned implicitly within a diffusion model.
@@ 3.3.1 Bayesian posterior
As before, we first use the Kalman filter formalism. Since the forecast is treated as an additional observation, we simply find a Kalman filter of the following form
@@
but with an extended observation and observation error covariance matrix, viz.
@@
Repeating the same calculation as in Section 3.1.1 gives the elements of the posterior mean and posterior variance:
@@
where $[\mathbf{f}_{k}]^{j}=f$ We note that as $r_{f}\to\infty$, that this posterior disregards the forecast and we recover the results from Section 3.3.1.3.1.1 as expected.
@@ 3.3.2 Diffusion DA
In this case, the typical diffusion model loss function would be modified to draw its training set from the extended joint posterior, i.e. ,
$p(\mathbf{x}_{k},\mathbf{y}_{k},\mathbf{f}_{k})=p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{f}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k})$ (64)
where we emphasize that the prior here is the climatological one. This implies the following loss
$\ell=E_{\mathbf{x}_{k},\mathbf{y}_{k},\mathbf{f}_{k}\sim p(\mathbf{x}_{k},\mathbf{y}_{k},\mathbf{f}_{k})}E_{n\sim\mathcal{N}(0,t^{2})}||\mathbf{x}_{k}-\mathbf{D}(\mathbf{v}=\mathbf{x}_{k}+n,t,\mathbf{y}_{k},\mathbf{f}_{k})||_{2}^{2},$ (65)
The minimum of (65) for a linear, Gaussian problem is as before a kind of Kalman filter of the form
$\mathbf{D}(\mathbf{v},t,\mathbf{y}_{k},\mathbf{f}_{k})=E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k},\mathbf{f}_{k})}[\mathbf{x}_{k}]=\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}$ (66)
where
$\mathbf{R}=\begin{bmatrix}t^{2}\mathbf{I}_{n_{x}}&\mathbf{0}_{n_{x}\times n_{o}}&\mathbf{0}_{n_{x}}\\ \mathbf{0}_{n_{o}\times n_{x}}&r\mathbf{I}_{n_{o}}&\mathbf{0}_{n_{o}\times n_{x}}\\ \mathbf{0}_{n_{x}}&\mathbf{0}_{n_{x}\times n_{o}}&r_{f}\mathbf{I}\end{bmatrix},\quad\hat{\mathbf{H}}=\begin{bmatrix}\mathbf{I}_{n_{x}}\\ \mathbf{H}\\ a\mathbf{I}_{n_{x}}\end{bmatrix},\quad\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}\\ \mathbf{y}_{k}\\ \mathbf{f}_{k}\end{bmatrix}.$ (67)
We then use this denoiser to determine the associated diffusion model, viz.
$d\mathbf{v}=-\frac{2}{t}\Big{[}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\hat{\mathbf{H}}^{T}+\mathbf{R}]^{-1}\hat{\mathbf{y}}-\mathbf{v}\Big{]}dt+\sqrt{2t}d\overline{\beta}_{t}$ (68)
Upon repeating the same steps in Section 3.3.1.3.1.2 we find that the mean of the observed variable is
@@
which, for $T\to\infty$ is equal to $\gamma$ which is the variance computed via the Kalman formalism in (63).

### v2

@@ 4.3.1 Bayesian posterior
The Bayesian posterior of interest here is (1). We compute the posterior means and variances but in the interest of clarity we focus on the steady state, i.e., $\mathbf{P}^{f}\to\mathbf{P}^{f}_{\infty}$, $\mathbf{P}_{a}\to\mathbf{P}^{a}_{\infty}$, $\mathbf{K}\to\mathbf{K}_{\infty}$ for large $k$ and, therefore, we have that
@@
$\mathbf{K}_{\infty}$ $=\mathbf{P}^{f}_{\infty}\mathbf{H}^{T}\left(\mathbf{H}\mathbf{P}^{f}_{\infty}\mathbf{H}^{T}+r\mathbf{I}\right)^{-1},$ (61)
Because both the cycling prior and the observation likelihood are Gaussian for this problem these equations can be derived, like in the previous section, by completing the square or, for those familiar with the Kalman formalism, by comparison to the Kalman filter. The steady state mentioned above is reached in linear, Gaussian systems when the observation operator, $\mathbf{H}$, and the observation error covariance, $\mathbf{R}$, matrices are constants in time.
The posterior mean does not converge to a steady state, but at a large time $k$ can be written as
@@
where $\overline{\mathbf{x}}_{fk}$ is the forecast mean, i.e., the posterior mean of the previous cycle evolved forward by the model (23). Because we observe $n_{y}$ elements of $\mathbf{x}_{k}$ directly, all matrices above are diagonal and it is sufficient to focus on one observed variable of the state vector. An element of the Kalman gain corresponding to an observed system state is
$\mathbf{K}_{\infty}]_{j}=\frac{\alpha^{j}}{\alpha^{j}+r},$ (63)
which can be obtained from (61) by setting $\mathbf{H}\to 1$, $\mathbf{P}_{\infty}^{f}\to\alpha^{j}$ and $\mathbf{I}\to 1$. Similarly, we obtain the posterior mean and variance of an observed state variable as
$[\overline{\mathbf{x}}^{a}_{k}]^{j}$ $=[\overline{\mathbf{x}}_{fk}]^{j}+\frac{\alpha^{j}}{\alpha^{j}+r}([y_{k}]^{j}-[\overline{\mathbf{x}}_{fk}]^{j}),$ (64)
@@
Because the data carry no information about unobserved grid points in our simplified setup, unobserved state variables are simply random draws from the cycling prior.
@@ 4.3.2 Diffusion DA
We consider a diffusion DA system that is retrained at every cycle using a time evolving training dataset. We will show that there is a way to construct such a diffusion DA system that samples this Bayesian posterior distribution with a cycling prior, i.e., the posterior distribution typically targeted by ensemble DA (see Bao et al. (2024) for another example of such a diffusion DA system).
@@
    We use this diffusion model to randomly sample $p(\mathbf{x}_{0}|\mathbf{y}_{0})$ to generate a large training set.
@@
    We use the forecast model, which can be either a numerical solution to (23) or ML/AI based, to push each member of this set forward to the time of the next set of observations, $\mathbf{y}_{1}$.
@@
    We then use these forecasts as the training dataset representing the cycling prior $p(\mathbf{x}_{1}|\mathbf{y}_{0})$ to build a new diffusion model that produces samples from $p(\mathbf{x}_{1}|\mathbf{y}_{1},\mathbf{y}_{0})\propto p(\mathbf{y}_{1}|\mathbf{x}_{1})p(\mathbf{x}_{1}|\mathbf{y}_{0})$ using the same methodology as in Section 4.4.2.4.2.2.
@@
    We then repeat steps 3 and 4 for each new cycle.
We imagine we have done this up to the $k^{th}$ cycle and that the covariances have converged to their steady-state values – which they will, assuming the diffusion is trained with sufficiently many samples.
The reverse process is
$d\mathbf{v}=-2t\nabla\log(p_{t}(\mathbf{v}|\mathbf{y}_{k},\dots,\mathbf{y}_{1}))\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t},$ (66)
and we use Tweedie’s formula to compute the score
$\nabla\log p_{t}(\mathbf{v}|\mathbf{y}_{k},\dots\mathbf{y}_{1})=\frac{E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k},\dots,\mathbf{y}_{1})}[\mathbf{x}_{k}]-\mathbf{v}}{t^{2}}.$ (67)
As before, we avoid neural networks and compute the conditional mean analytically, repeating the steps of Section 4.4.2.4.2.2, but replacing the climatological prior in (35) with the cycling prior $p(\mathbf{x}_{k}|\mathbf{y}_{k-1},\dots,\mathbf{y}_{1})$, i.e., the time-evolving (regenerated at each cycle) training dataset. The result is
$E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{y}_{k},\mathbf{y}_{k-1},...)}[\mathbf{x}_{k}]=\overline{\mathbf{x}}_{fk}+\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}+\hat{\mathbf{R}}]^{-1}\hat{\mathbf{y}}$ (68)
with $\hat{\mathbf{H}}$ and $\hat{\mathbf{R}}$ as before, but where we now have
$\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}-\overline{\mathbf{x}}_{fk}\\ \mathbf{y}-\mathbf{H}\overline{\mathbf{x}}_{fk}\end{bmatrix}$, (69)
due to the cycling prior. Plugging the result into the reverse process yields
$\text{d}\mathbf{v}=-\frac{2}{t}\left[\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}[\hat{\mathbf{H}}\mathbf{P}^{f}_{\infty}\hat{\mathbf{H}}^{T}+\hat{\mathbf{R}}]^{-1}\hat{\mathbf{y}}-(\mathbf{v}-\overline{\mathbf{x}}_{fk})\right]\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t},$ (70)
Due to our simple observation operator, all matrices above are diagonal and we can compute the reverse process element-wise to be
@@
Note that $v(0)$ is the Gaussian random variable of interest and we compute its mean
$\langle v(0)\rangle=\overline{x}+\frac{\frac{\alpha^{j}r}{\alpha^{j}+r}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2}}(\langle v(T)\rangle-\overline{x})+\frac{(\alpha^{j})^{2}r}{(\alpha^{j}+r)^{2}}\frac{T^{2}}{\frac{\alpha^{j}r}{\alpha^{j}+r}(\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2})}(y-\overline{x}),$ (73)
which, in the limit as $T\rightarrow\infty$, simplifies to
$\langle v(0)\rangle_{T\rightarrow\infty}=\overline{x}+\frac{\alpha^{j}}{\alpha^{j}+r}(y-\overline{x}).$ (74)
The above result is equal to the cycling posterior mean derived from the Bayesian posterior in (64). Next, we compute the variance of $v(0)$
@@
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{(\frac{\alpha^{j}r}{\alpha^{j}+r})^{2}}{(\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2})^{2}}\langle(v(T)-\langle v(T)\rangle)^{2}\rangle+\Bigg{\langle}\Bigg{(}\frac{\alpha^{j}r}{\alpha^{j}+r}\int_{T}^{0}\frac{\sqrt{2t}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+t^{2}}\text{d}\overline{\beta}_{t}\Bigg{)}^{2}\Bigg{\rangle}\end{split}$ (75)
[…]
$\begin{split}\langle(v(0)-\langle v(0)\rangle)^{2}\rangle=\frac{\frac{(\alpha^{j})^{2}r^{2}}{(\alpha^{j}+r)^{2}}T^{2}}{(\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2})^{2}}+\frac{\alpha^{j}r}{\alpha^{j}+r}\frac{T^{2}}{\frac{\alpha^{j}r}{\alpha^{j}+r}+T^{2}}\end{split},$ (76)
and in the limit $T\rightarrow\infty$ we find
$\langle(v(0)-\langle v(0)\rangle)^{2}\rangle_{T\rightarrow\infty}=\frac{\alpha^{j}r}{\alpha^{j}+r},$ (77)
which is the posterior variance we obtained the Bayesian posterior in (65).
We have thus shown that a cycling prior corresponds to a time-evolving (regenerated at each cycle) training dataset for diffusion DA. Put differently, a diffusion DA system that targets the same Bayesian posterior distribution as an ensemble DA requires re-training the denoiser at each cycle.
4.4 Diffusion DA with an extended likelihood
We now consider diffusion DA systems that ingest observations and a forecast, denoted by $\mathbf{f}_{k}$. The forecast propagates information from the past into the current DA cycle and we assume that it is derived from the posterior mean at the previous cycle k1k-1. We subsequently treat the forecast as a “pseudo” observation, which emulates one of many possible methods of supplementing a diffusion DA system with a forecast. We make our assumptions based on what we think is most reasonable at the time of writing.
In terms of priors, likelihoods and posterior distributions, we now consider the state being conditioned on the observations and the forecast $\mathbf{f}_{k}$, i.e., the Bayesian posterior distribution of interest is
$p(\mathbf{x}_{k}|\mathbf{y}_{k},\mathbf{f}_{k})\propto p(\mathbf{y}_{k}|\mathbf{x}_{k})p(\mathbf{f}_{k}|\mathbf{x}_{k})p(\mathbf{x}_{k}),$ (78)
where, $p(\mathbf{x}_{k})$ is the climatological prior, $p(\mathbf{y}_{k}|\mathbf{x}_{k})$ is the usual Gaussian observation likelihood and $p(\mathbf{f}_{k}|\mathbf{x}_{k})$ is an “extended” likelihood to include the forecast $\mathbf{f}_{k}$, which we assume is independent of the observation at time $k$. The extended likelihood requires that we create a model for how the forecast is related to the “truth,”. As we demonstrate in Appendix E, this relationship is linear for our linear/Gaussian test problem and we assume that
$\mathbf{f}_{k}=a\mathbf{x}_{k}+\bm{\varepsilon}_{f},\quad\bm{\varepsilon}_{f}\sim\mathcal{N}(\mathbf{0},r_{f}\mathbf{I}),$ (79)
where $0\leq a\leq 1$ and $r_{f}$ are scalars, which we determine later on.
@@ 4.4.1 Bayesian posterior
As before, one can either complete the square within equation (78) or make use of the Kalman filter formalism to compute the posterior mean and covariance explicitly. Since the forecast is treated as an additional observation, we find
@@
where
@@
Repeating the same calculation as before gives the elements of the posterior mean and posterior variance:
@@
where $\mathbf{f}_{k}]^{j}=f$. We note that as $r_{f}\to\infty$, that this posterior disregards the forecast and we recover the results from Section 4.4.2.4.2.1 as expected.
@@ 4.4.2 Diffusion DA
Given observations and a forecast, the reverse process of a diffusion model becomes
$d\mathbf{v}=-2t\nabla\log(p_{t}(\mathbf{v}|\mathbf{y}_{k},\mathbf{f}_{k}))\text{d}t+\sqrt{2t}\text{d}\overline{\beta}_{t},$ (84)
and we use Tweedie’s formula to compute the score
$\nabla\log p_{t}(\mathbf{v}|\mathbf{y}_{k},\mathbf{f}_{k})=\frac{E_{\mathbf{x}_{k}\sim p(\mathbf{x}_{k}|\mathbf{v},\mathbf{y}_{k},\mathbf{f}_{k})}[\mathbf{x}_{k}]-\mathbf{v}}{t^{2}}.$ (85)
We again avoid neural networks and compute the conditional mean analytically, repeating the steps of Section 4.4.2.4.2.2, but extending the likelihood by the forecast, so that
$\hat{\mathbf{y}}=\begin{bmatrix}\mathbf{v}\\ \mathbf{y}_{k}\\ \mathbf{f}_{k}\end{bmatrix},\quad\hat{\mathbf{H}}=\begin{bmatrix}\mathbf{I}_{n_{x}}\\ \mathbf{H}\\ a\mathbf{I}_{n_{x}}\end{bmatrix},\quad\hat{\mathbf{R}}=\begin{bmatrix}t^{2}\mathbf{I}_{n_{x}}&\mathbf{0}_{n_{x}\times n_{o}}&\mathbf{0}_{n_{x}}\\ \mathbf{0}_{n_{o}\times n_{x}}&r\mathbf{I}_{n_{o}}&\mathbf{0}_{n_{o}\times n_{x}}\\ \mathbf{0}_{n_{x}}&\mathbf{0}_{n_{x}\times n_{o}}&r_{f}\mathbf{I}\end{bmatrix}$. (86)
Repeating the same steps as in Section 4.4.2.4.2.2 but with $\hat{\mathbf{R}}$, $\hat{\mathbf{H}}$ and $\hat{\mathbf{y}}$ as above gives the the mean of the observed variable,
@@
which, for $T\to\infty$ is equal to $\gamma$ which is the variance of the Bayesian posterior in (83).